OpenStudy (anonymous):

4/5 + 5/6 ?? simplify

OpenStudy (anonymous):

Denominators are 5 and 6 Least common denominator is 30 We make both the denominators 30 4/5 = 4/5 * 6/6 = 24/30 5/6 = 5/6 *5/5 = 25/30 4/5 + 5/6 = 24/30 + 25/30 = (24 +25)/30 =49/30 = 1 19/30 ( 1 is the whole number and 19/30 is fraction)